OpenStudy (anonymous):

What am I doing wrong? (Probability) The question is: A large bag of marbles contains equal amounts of each of 5 colors (red, orange, yellow, green, and blue). Joaquin selects 1 marble, looks at it, puts it back, and then selects another. Find the probability of Joaquin selecting a yellow marble both times. ___ I figured I had to find the probability of picking a Y out of 10 marbles. (I just chose ten as it would be easiest?) So that means there would be two Y marbles in said bag. So, total = 10 And each time, you should have 2/10 as the probability of piking out a Y marble.

OpenStudy (anonymous):

2/10 + 2/10 = 4/10 = 2/5 = 0.4 But that is not an option?

OpenStudy (anonymous):

The choices are: A. 0.5 B. 0.04 C. 0.008 D. 0.25 My question is what am I doing wrong?

OpenStudy (acxbox22):

you can just take it out of 5 1 yellow out of 5 total for 1st time 1 yellow out of 5 total for second time now the mistake you did was too add them...

OpenStudy (anonymous):

OHH! Understood! Thank you @acxbox22 :)

OpenStudy (acxbox22):

:)